APPENDIX C: Sample U of O Thesis Pages from the “Grey Book”

This appendix consists of TEX output representing sample U of O thesis pages taken from the University of Oregon Style and Policy Manual for Theses and Dissertations, University of Oregon Graduate School, revised September 1987.

A machine-readable copy of the raw TEX commands used to typeset these pages is available to OREGON VAX users as:

      TEX$DISK:[TEX.SAMPLES]UOTHESIS.TEX

On other systems get a copy of that file via anonymous FTP.

FTP to DECOY.UOREGON.EDU (128.223.32.19) and see directory pub/tex/samples.

We’ve done our best to make these sample pages meet U of O Graduate School requirements as outlined in the “Grey Book,” but in the final analysis, you are the one who has to please the Graduate School and your committee, not us.

We strongly urge you to consult on an early and ongoing basis with the Graduate School editor and your committee members for authoritative assistance in meeting dissertation content and format requirements.

ANALYSIS OF UNUSUAL CONFORMATIONS OF OLIGIO

AND POLYNUCLEOTIDES AND PEPTIDES BY CLASSIC RAMAN SCATTERING

by GERALD ALAN THOMAS

A DISSERTATION Presented to the Department of Chemistry and the Graduate School of the University of Oregon in partial fulfillment of the requirements for the degree of Doctor of Philosophy August 1984 APPROVED:

Dr. Harry S. Truman

iii

\begin{matrix}c \\ ◯ \end{matrix} 1984 Jane Doe

An Abstract of the Dissertation of

Walter James Teasdale  for the degree of  Doctor of Philosophy
in the Department of Economics  to be taken  August 1984
Title: ADOPTION AND IMPLEMENTATION OF THE NORTHWEST POWER

PLANNING COUNCIL’S MODEL CONSERVATION STANDARDS

Approved:

Dr. Harry S. Truman

The Pacific Northwest Electric Power Planning and Conservation Act of 1980 and the Northwest Conservation and Power Plan are models for economic planning. It is an essential part of the Council’s resource portfolio. The Standards set energy use budgets for new buildings and (text continues)

VITA

NAME OF AUTHOR: Jane Ann Doe PLACE OF BIRTH: Portland, Oregon DATE OF BIRTH: January 1, 1950

GRADUATE AND UNDERGRADUATE SCHOOLS ATTENDED: University of Oregon San Francisco State College San Jose State College

DEGREES AWARDED:

Doctor of Philosophy, 1984, University of Oregon Master of Arts, 1972, San Francisco State College Bachelor of Arts, 1970, San Jose State College

AREAS OF SPECIAL INTEREST:

Psychology of Perception

PROFESSIONAL EXPERIENCE:

Research Assistant, Department of Psychology, Univeristy of

Oregon, Eugene, 1973-75 Teaching Assistant, Department of Psychology, University of

Oregon, Eugene, 1972-73

AWARDS AND HONORS:

Danforth Fellowship, 1976-77

PUBLICATIONS:

(if any publications, list here, using same style sheet you use for your bibliography; if none, omit this heading)

ACKNOWLEDGMENTS

The author expresses sincere appreciation to Professors Smith and Jones for their assistance in the preparation of this manuscript. In addition, special thanks are due to Mr. John Doe, whose familiarity with the needs and ideas of the community were helpful during the early programming phase of this undertaking. I also thank the members of the community councils for their valuable input. The investigation was supported in part by a Public Health Service Predoctoral Fellowship, Number 4 FL MH-21, 824-02, from the National Institutes of Mental Health, and by a grant from the National Science Foundation, SPI 801-2372, to Dr. Philip Smith of the University of Oregon.

DEDICATION

This manuscript is dedicated to my wife and children.

TABLE OF CONTENTS

Chapter  Page
I. INTRODUCTION .................................................  1
II. REVIEW OF LITERATURE .....................................  9
Current Research on Marital Status and Income ..............  9
Research on Sex Roles ..........................................  9

The Relationship Between Marital Status and

Income Following the Birth of the First Child ..............  26
Summary ........................................................  34
III. METHODOLOGY ................................................  41
Subjects .........................................................  42
Instruments ......................................................  44
General Procedures .............................................  53
IV. RESULTS ..........................................................  55
Reality Hypothesis Result ......................................  56
Tangent Hypothesis Result .....................................  57
Circular Hypothesis Result .....................................  59
Summary ........................................................  62
V. SUMMARY OF FINDINGS AND CONCLUSIONS .............  67
Major Findings ..................................................  68
Study Limitations ...............................................  71
Suggestions for Future Studies .................................  81

APPENDIX

A. SUBJECT PROFILES ............................................  90
B. INTERVIEW QUESTIONNAIRE ................................  98
C. LETTER TO PARTICIPANTS ..................................  99
D. OTHER DATA ....................................................  100
BIBLIOGRAPHY  .........................................................  123

LIST OF TABLES

Table  Page
1. Frequencies and Percentages for Geographic Data

of Subjects’ Marital Status and Average Age

as Categorized by Group and Income ..........................  77
2. Frequencies and Percentages for Geographic Data

of Subjects’ Sex, Marital Status, and Average

Age Distributed Across Operational Systems ..................  78
3. Mean, Standard Deviation, and Standard Error Scores
of Responses Integrating Input Compatibility .................  79
4. Summary Table for Frequencies and Percentages of
Responses Integrating Input Compatibility ....................  80
5. Summary Table for Marital Status Contrasted
with Sequential Output Data ...................................  81
6. Summary Table for One-Way ANOVA Performed on
Multiple Interactive Compatibility Data .......................  82
7. Summary Table for One-Way ANOVA Performed on
Functional Capabilities Data ...................................  83

CHAPTER I

INTRODUCTION

The spirit of this style manual is an inquiry into the parameters set by necessity in order to standardize publications from graduate students at the University of Oregon. The \underline{Manual} is to be used as a guide. It does not contain all requirements. Other style manuals should be considered.

Style Manuals

Other manuals are available at the University Bookstore for information. Since the student has to make a choice concerning style of the manuscript at an early date, we suggest using the library before deciding which style manual to purchase. The Graduate School suggests using the manuals published by the University of Chicago.

Chicago Style Manuals

There are two manuals available from the University of Chicago which the Graduate School uses in making final determination on questions not addressed in the Graduate School Style Manual.

\underline{Turabian}

The \underline{Manual for Writers of Term Papers, Theses, and Dissertations} (Turabian, 1973) is one resource for the student writing a thesis.

Assistant Administrator Nursing Service Director Specialty Units Surpervisor Nurse Unit Manager

Nurse  Charge Nurse  Nurse
Facilitator  Facilitator
7:00 A.M. to  3:00 P.M. to  11:00 P.M. to
3:00 P.M.  11:30 P.M.  7:30 A.M.
Registered  Registered  Registered
Nurse  Nurse  Nurse
Unit Secretary  Unit Secretary
8:00 A.M. to  4:00 P.M. to
4:00 P.M.  8:00 P.M.

FIGURE 4. Nursing services authority structure for the neonatal intensive care unit.

TABLE 12. Analysis of Variance of Mean Initial After-Effect for the Sighted Children, Ages 5 Through 16

Source of  Sum of  Mean
Variation  Squares  df  Square  F  p.
Between Groups  2,222.22  11  202.02  1.31 N.S.
Within Groups  19,822.12 129  153.55
Total  22,044.34 140

TABLE 13. Methods Teachers Say They Would Employ Under New Conditions as They Pertain to Increased Use of a Variety of Learning Activities and Projects

205  Rank
Teachers  Order of
Item Reason  Reporting  Response
No. %  Order
1.  More individual and small-group
projects 17 8.0 8
2.  More activities with practical

application to subject matter

being taught  13 6.1  15
3.  More creative activities  12 5.6  16
4.  More freedom of choice of
activities by pupils 7 3.3 33
5.  Enrichment activities for all
pupils in all areas of instruction  7 3.3  34
6.  More child-centered project work  6 2.8  38
7.  More use of field trips  3 1.4  70

5

APPENDIX

6

BIBLIOGRAPHY